OpenStudy (anonymous):

Find the focus of a parabola with a = 150 1/4. so a = 1/4c 150 1/4 = 1/4c I need help solving this for c.

OpenStudy (anonymous):

Ok well \[\Large \frac{1}{4}c=\frac{c}{4} \]

OpenStudy (anonymous):

And the opposite operation of division is ?

OpenStudy (anonymous):

multiplication. so i ignore the 150 for now?

OpenStudy (anonymous):

Well no. Is the 150 1/4 actually \[\Large 150\frac{1}{4}~or~150.25?\]

OpenStudy (anonymous):

Like is that what number it is or is it just hte question number, question #150

OpenStudy (anonymous):

150 1/4

OpenStudy (anonymous):

Ok well you gotta convert it to an improper fraction, do you know how to do that?

OpenStudy (anonymous):

150 times 4 add 1

OpenStudy (anonymous):

Ok, and that gives you \[\Large \frac{601}{4}\]

OpenStudy (anonymous):

Now you're gonna multiply 4 by both sides to get c by itself

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

601 = 1/c

OpenStudy (anonymous):

Remember it was c/4 so multiplying both sides by 4 would have a 4 in the numerator and denominator and they cancel out adn you're left with c=601

OpenStudy (anonymous):

okay, thank you!!

OpenStudy (anonymous):

You're welcome